OpenStudy (anonymous):

completely factor the trinomial or indicate if its not factorable: 15p^4+96p^3+105p^2

OpenStudy (anonymous):

3p^2(5p^2+32p+35) d=32^2-4x5x35=324>0 it is factorable

OpenStudy (anonymous):

\[3p ^{2}(p+5)(5p+7)\] p=0,-5,-1.4